No tunnel works today.
- Detach the remaining white tapes just before closing the chamber; be noted that there are white tapes on the PEEK connectors at the inside of the flanges.
- Torque completely the bolts which are fixing the upper and the lower base plates.
- Insert metal gaskets before torquing the ICF203 flanges with Dsub connectors.
- Attach an oplev viewport
- Setup the oplev for TMS-VIS
- Consider why the green ghost beam is not observed; that was observed in TMSY in this April.
- Modify MEDM screen
- Add DC PD and so on.
Cable Dsub 15pin for TMS-VIS OpLev.
- Return plastic boxes.
- Install rails in digital rack.
- They may be necessary under LVDT distributor, geophone distributor, and WHT board.
- Nagano is not sure it is necessary under HP coil driver.
- Measure the length between BRT lens3 and GLNS1
- Black tubes attachable to the DC-PDs or seetabale in front of the DC-QPDs for light shielding.
- Screws to screw Dsub 15pin connectors to QPD amplifiers
- Already screwed for IR and GR QPDs. Not yet for TMSX VIS oplev.
- Nice way to turn off the ceiling LED light
- Covers for cables running on the floor
- Shield box against environmental light
- Wind shielding... (it would be somewhat difficult as two beams are running in the different directions from the viewport window)